OpenStudy (anonymous):

Medals!!!! The table below shows the outputs y for different inputs x: Input (x) 1 1 3 3 Output (y) 4 2 12 6 Part A: Does the table represent y as a function of x? Justify your answer. Part B: The total cost f(x), in dollars, for renting a paddleboat for x hours is shown below: f(x) = 20 + 10x What is the value of f(120), and what does f(120) represent?

OpenStudy (anonymous):

For something to be a function, an x value (input) must not be repeated. We cannot input 1 and get both 4 and 2 as outputs. Dont forget however, we /CAN/ have same outputs for different inputs, e.g Input = {1,2,3,4} Output = {0,0,0,0} IS VALID - We got same outputs for unique inputs. Input = {1,2,3,3} Output = {12,2,32,13} IS INVALID - We repeated inputs and got different outputs. Part B: f(x) = 20 +10x. Where x is the number of hours. f(x) represents the cost. In essence this function tells you how much it costs to rent a boat out for some number of hours. x is hours. f(x) is cost. So f(120) = Cost of boating for 120 hours.(lots of boating) To compute the value of f(120) simply plug in 120 for x in the equation given, f(120) = 20 + 10(120) f(120) = 20+1200 f(120) = 1220 dollars.